OpenStudy (blackfireskull):

solve 3x-1=10 A) 3 B) 11/3 C) 3/11

OpenStudy (blackfireskull):

are you sure? .-.

OpenStudy (blackfireskull):

i'm so confused! >.< @Abhisar could you help?

OpenStudy (abhisar):

3x-1 = 10 First you add 1 on both sides 3x -1 =10 + 1 What do you get?

OpenStudy (blackfireskull):

3x=11

OpenStudy (abhisar):

Yes, finally you divide both sides with 3

OpenStudy (anonymous):

I am sorry for confusing you. my sister was screaming in my ear

OpenStudy (blackfireskull):

so it is 11/3?

OpenStudy (abhisar):

Correct!

OpenStudy (blackfireskull):

could you help with one more? @Abhisar

OpenStudy (abhisar):

Sure, open a new thread.